OpenStudy (anonymous):

What are the solutions of 3x2 + 4x = –5?

OpenStudy (anonymous):

Please Help!

OpenStudy (cwrw238):

3x^2 + 4x + 5 = 0 you can solve this by using quadratic formula for ax^2 + bx + c = 0 x = ( (-b +-sqrt(b^2 - 4ac)) / 2a

OpenStudy (cwrw238):

here a = 3 . b = 4 and c = 5 plug these into the above formula

OpenStudy (anonymous):

I still dont get it, @Muskan can you help?

OpenStudy (anonymous):

3x2+4x+5=0 -4+-\[\sqrt{4-4*3*5/2*3}\] -4+-\[\sqrt{-56/6}\]

OpenStudy (anonymous):

is it -66?

OpenStudy (anonymous):